In photoelectric effects the work function of a metal is 3.5 V. The emitted electrons can be stopped by applying a potential of -1.2 V. Then


(a)The energy of the incident photon is 4.7 eV


(b)The energy of the incident photon is 2.3 eV


(c)If higher frequency photon be used, the photoelectric current will rise.


(d)When the energy of photon is 3.5 eV, the photoelectric current will be maximum


Grade:upto college level

1 Answers

Amit Saxena
10 years ago

(a)

hv= W0 + Ek = 3.5 + 1.2 = 4.7 eV